• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E. Cisbani

    Log entry time 08:00:35 on June28,2005

    Entry number 147322

    keyword=Owl Shift summary

    Owl Shift summary 06/28

    01:00-03:30 Drifting of the waterfall target tacheo speed for 600 to 615 approx.

    01:30-2:23 The Halla:p energy fluctuated by 200 keV (instead of few tens of keV); MCC fixed it tuning the fast feedback. Run affected 4930 and part of 4931. Beam position drifting in run 4931 probably related to the fixing of the energy oscillations.

    04:15 RF-separator problem, beam off for the rest of the shift

    Runs Summary

    previous shift-00:29 Run 4928 (H2O speed = 603.4)

    00:30-01:28 Run 4929 (H2O speed = 605.2)

    01:29-02:19 Run 4930 (H2O speed = 601.8) possible energy fluctuation (see entry 147308

    02:20-02:49 Run 4931 (H2O speed = 605.9) beam position on target unstable (see entry 147309, run manually stopped at about 500 k events

    02:50-03:49 Run 4932 (H2O speed = 609.7)

    03:50-04:30 Run 4933 (H2O speed = 610.8), run manually stop, beam off for 0.5 hour